### v4

# Anomalous diffusion in convergence to effective ergodicity

Abstract
The nature of diffusion is usually studied for particles or dynamics generating trajectories over time. Similar in principle, these studies can be executed in tracking how a given function of observable properties evolve over time akin to a system’s particle motion, so-called functional-diffusion. This is not the same as systems’ own trajectories but can be considered as a meta-trajectory. Following this idea, we measure how an approach to ergodicity evolves over time for the observed magnetization of a full Ising model with an external field. We compute the functional’s diffusive behavior depending on a range of temperatures via Metropolis and Glauber single-spin-flip dynamics. System’s ensemble-averaged dynamics are computed using expressions from the exact solution. Power-laws on the approach to ergodicity provide the classification of anomalies in the functional-diffusion, demonstrating non-linear anomalous behavior over different temperature and field ranges.
keywords:
ergodicity, power-laws, Ising models, lattice dynamics, monte carlo, functional-diffusion, Glauber dynamics, Metropolis dynamics
articletype: Paper
1 Introduction
Brownian motion is probably one of the landmark concepts in statistical physics attracted Einstein’s interest early on [1]. Its importance in formulating statistical mechanics recently reviewed within its 250th anniversary publications [2]. Major observable in tracing Brownian motion appears as how accumulated displacement curve behaves over time, i.e., linear relationship without intercept corresponding to normal diffusion. If the displacement curve of trajectories shows a power-law scaling over time other than the linear relationship, then the behaviour is called anomalous diffusion [3, 4, 5]. There is a recent surge in interest in using machine learning techniques in the analysis of anomalous diffusion data [6, 7, 8, 9], for example, the characterisation using transformer architectures is quite novel given the larger interests in attention mechanism in machine learning [10].
Along these lines, the dynamics of cooperation among assembly of independent units has been studied in this context [11], such as model of magnetic material [12, 13, 14] and the state of a neuron [15, 16]. Measuring the diffusion behaviour in this discrete cases would not only yields to mathematically challenging consequences, provides insights into characteristics of the Ising model. In our described scenario does not apply to trajectories, rather a functional of trajectories. This distinction is the core driver in our work that a functional on the function that explains convergence to ergodicity for total magnetisation is the diffusing, rather than the trajectory of the discrete units of the Ising model.
Convergence to ergodicity in this kind of cooperation dynamics is explored and established [17] using Thirumalai-Mountain (TM) fluctuation metric [18, 19] that is adapted for the Ising model’s total magnetization. In this study, we investigate the power-laws that emerged from the time evolution of the rate of effective ergodic convergence under the parameters that give rise to strong correlations and anomalous diffusion in convergence. Understanding ergodicity for these circumstances is not only interesting due to fundamental importance in statistical mechanics [20] and anomalous diffusion in general but also for its implications in real-world applications, such as understanding disruption in neural networks for dementia [21], realization of associative memory in a solid-state device [22], nature of economic utility [23] and optical lattice dynamics [24].
The formulation of the Ising model is reviewed in Sec. 2. In Sec. 3, the concept of effective ergodicity is explained. Measuring convergence to ergodicity is shown in Sec. 4; this corresponds to the functional-diffusion concept we introduced. Computation of power-law will be shown in Sec. 5 and our experimental results in Sec 6. The conclusion is given in Sec. 7.
2 Ising-Lenz System
A lattice with $N$ sites, labeled as $\{s_{i}\}_{i=1}^{N}$ which can take two values, such as $\{1,-1\}$. These values imply spin up or down in the Ising Model [12, 13, 14] or an activation in neuronal systems [15, 16]. The total energy, the Hamiltonian of the system, can be written with two interactions, one due to nearest-neighbors (NN) and one due to an external field, with coefficients $J$ and $H$ respectively,
$\mathcal{H}({s_{i}}{i=1}^{N},J,H)=J\Big((\sum_{i=1}^{N-1}s_{i}s_{i+1})+(s_{1}s_{N})\Big)+H\sum_{1}^{N}s_{i}.$ (1)
The term $s_{1}s_{N}$ is imposed by the periodic boundary conditions, which provide translational invariance, making the model a closed chain of interacting units. The thermal scale is expressed with $\beta=\frac{1}{k_{B}T}$ and the corresponding coefficients for NN and external field are scaled by this,
$K=\beta J,\qquad h=\beta H.$ (2)
The analytic expression for finite size magnetisation $M_{E}(N,\beta,H)$ is obtained by using the Transfer Matrix method. Whereas the time evolution of $M_{E}(N,\beta,H)$ i is computed via the Monte Carlo procedure with Metropolis and Glauber single-spin-flip dynamics. More details can be found in the previous work [17].
Figure 1: Diagnostic plots of $40$ different runs for $N=1024$ with Glauber dynamics for the evolution of $K(t)$ log-log regression (a) A field at $H=1.0$ and inverse temperature $\beta=1.16$ (b) A field at $H=1.4$ and inverse temperature $\beta=1.0$. We identify optimal time starting fit with minimal Kolmogorov-Simirnov statistic on a grid-search.
Figure 2: Uncertainty quantified two power-laws’s $\alpha$ exponents over temperature ranges (a) the time-evolution of $K(t)$, via log-log regressions. (b) the distribution of $\Gamma(t)$ via analytical expressions from Newman et. al.. Here we plot absolute value as $\alpha$. In the fit, functional form is negative $\alpha$.
3 Ergodic Dynamics
A form of ergodic dynamics suggested by Boltzmann that trajectories of a many-body system will reach phase-space partitions where it’s likely to be in a thermodynamic equilibrium [20]. At this point, ensemble averages and time averages of the system produce close to identical measures. This implies for a given observable $g$ over a fixed phase-space point, ensemble-averaged value can be recovered by time-averaged values, $g(t)$ from $t_{0}$ to $t_{N}$,
$\langle g\rangle=lim_{t_{N}\to\infty}\int_{t_{0}}^{t_{N}}g(t)dt,$ (3)
where $\langle\rangle$ indicates ensemble-averaged value. This kind of basic definition is not standard in the literature [25, 18, 26]. Other forms of ergodicity demand that system should visit all available phase-space partitions, which might not be possible for a finite physical system, and moreover feasibility of this type of ergodicity is questioned [27]. In practical terms, since partitions of the phase-space are clustered around few regions, effective ergodicity can be attained quickly [18].
On the other hand, ergodicity for the single-spin-flip dynamics, essentially a Markov process, defined as accessibility of any given state point to another state point over time [26, 28]. In this sense, Monte Carlo procedure used in this work is ergodic by construction for long enough times.
4 Ergodicity convergence as functional-diffusion
Effective ergodic convergence, $\Omega_{G}(t)$ for a given observable $g$, can be constructed based on the fact that identical parts of cooperating units have identical characteristics in thermodynamic equilibrium [18]. This is realized by the Thirumalai-Mountain (TM) $G$-fluctuating metric [18, 19]. Applications of ergodicity with TM metric is quite widely used also with recent interests, biomolecular simulations [29], physical chemistry and machine learning landscapes [30], seismology [31, 32, 33, 34, 35, 36], neuromorphic computing [37, 38] and artificial spin ice [39, 40].
TM metric at a given time $t_{k}$ reads,
$\Omega_{G}(t_{k})=\frac{1}{N}\sum_{j=1}^{N}\big[g_{j}(t_{k})-\langle g(t_{k})\rangle\big]^{2},$ (4)
Functional-diffusion occurs for the development of a function of an observable over time, $F[O(t)]$, measured as its displacement from the initial conditions.In this sense, time-evolution of inverse of $\Omega_{G}(t_{k})$ can be considered as it is in functional-diffusion
where time-averaged per unit expressed as $g_{j}(t_{k})$ and $\langle g(t_{k})\rangle$ is the instantaneous ensemble-averaged over all units,
$g_{j}(t_{k})$ $=$ $\frac{1}{k}\sum_{i=0}^{k}g_{j}(t_{i}),$ (5)
$\langle g(t_{k})\rangle$ $=$ $\frac{1}{N}\sum_{j=1}^{N}g_{j}(t_{k}).$ (6)
Consequently, the rate of ergodic convergence is measured as
$\Gamma_{G}(t)=\frac{\Omega_{G}(t)}{\Omega_{G}(0)}\to\frac{1}{tD_{G}},$ (7)
where diffusion coefficient expressed as $D_{G}$. A similar measure of ergodicity is used in simple liquids [18, 41], and earthquake fault networks [42, 43]. The inverse rate of ergodic convergence can also be defined by $K_{G}(t)$,
$K_{G}(t)=\frac{\Omega_{G}(0)}{\Omega_{G}(t)}\to t\cdot D_{G}.$ (8)
Using $\Omega_{G}$ for the total magnetization in Ising Model at time $t_{k}$ as a function of temperature and external field values is expressed as
$\Omega_{M}(t_{k},N,\beta,h)$ $=$ $\big[M_{T}(t_{k})-M_{E}\big]^{2},$ (9)
$M_{T}$ $=$ $\frac{1}{k}\sum_{i=0}^{k}M(t_{i}),$
where $M_{T}(N,\beta,h)$ and $M_{E}(N,\beta,h)$ correspond to time-averaged and ensemble-averaged total magnetization. Since, ensemble averaged is computed analytically, this makes our approach using a modified TM-metric, leading to more accurate assessment of the ensemble averaging. Note that exact expression for $M_{E}$ is used [17].
5 Functional-diffusion: Two power laws
We would like to investigate the behaviour of the functional-diffusion in two power laws. The first-one is similar to the one investigated for displacement vector from the literature, but over ergodicity convergence function $K$ we defined above,
$K_{G}(t)\to C\cdot t^{\alpha}.$ (11)
whereby $C$ is the generalized diffusion coefficient, $t$ is the Monte Carlo time and for the positive $\alpha$ exponent. We call this time power-laws. The second type of power-law we seek is over the distribution of $\Gamma(t)$,
$P(\Gamma_{G}(t))\to\Gamma_{G}(t)^{-\alpha}.$ (12)
This is identified as distributional power-laws, whereby this is related to Lèvy flights or jump distributions [4], we consider cumulative jumps of approach to ergodicity here.
We will explicitly state which power-laws we are working on, not to confuse $\alpha$-s. The power laws in complex systems and computing them on a given empirical data are studied in depth using techniques pioneered by Newman et. al. [44, 45]. We did an extensive bias corrected bootstrapping for determining uncertainties on scaling exponents, diffusion coefficient, fit diagnostics, KS-distances (Kolmogorov-Simirnov Statistic) and adjusted R-squared, and autocorrelation times. KS-distance was critical in determining where to start to perform log-log regression depending on the parameters, this is done in an automated way with finding minimum KS-statistic over a search grid of time values. Visual inspection diagnostic is done for every parameters and diagnostic visualisations are developed and automated.
We also computed autocorrelations times using time self-correlations of averaged magnetisation, via location of the plateau of the correlation times $C(t)$ of averaged-magnetisations,
$C(t_{k})=\frac{1}{t_{k}}\sum_{i=0}^{k}M(t_{0})M(t_{i})$
. We identify that relaxation changes in the regions where anomalous convergence appears.
A Finite Size-Scaling (FSS) analysis could show a universality of the results depending on the number of discrete units and temperatures for the Ising model. We made the Following FSS assumptions, for temperature and size dependence
$u(\beta)=(\beta-\beta_{c})N^{a}$
scaling function $f(u)=A+Bu^{c}$, then power-law exponents $\alpha$ with the critical temperature $\beta_{c}$,
$\alpha(T,N)=N^{b}f(u)$
We identified the FSS exponents $a$, $b$, $c$ and coefficients $A$, $B$ with nonlinear optimization on the power-laws over temperatures for time power-laws.
Figure 3: Uncertainty quantified two power-laws’s $\alpha$ exponents over field ranges at $\beta=1.0$ (a) the time-evolution of $K(t)$, via log-log regressions. (b) the distribution of $\Gamma(t)$ via analytical expressions from Newman et. al.. Here we plot absolute value as $\alpha$. In the fit, functional form is negative $\alpha$.
Figure 4: Uncertainty quantified power-laws of the time-evolution of $K(t)$, via log-log regressions, diffusion coefficients (a) Over temperature ranges. (b) Over field ranges.
6 Experimental Results
We run extensive Metropolis and Glauber dynamics for the full Ising model and measure the rate of ergodic convergence and its inverse, $\Gamma(t)$ and its inverse $K(t)$. Here we present the result with our findings.
An of evolution of $K(t)$ is shown for 40 independent runs in Figure 1a and 1b for different temperature and field conditions, with it’s diagnostic of where we start fitting via optimal KS distance analysis. These plots served as a visualisal diagnostics in ensuring each temperature, size or field cases are inspected that the fitted log-log regression represents indeed a power-law region within the simulated MC time, whereby we record $\Gamma(t)$ are acceptance times.
Figure 5: Uncertainty quantified autocorrelation times over $200$ runs. (a) A diagnostic plot for identification of autocorrelation time. (b) Autocorrelation times over temperature ranges and three different sizes.
ikBT alphaMean alphaLower alphaUpper adjR2Mean adjR2Lower adjR2Upper
$0.50$ $0.0179$ $0.0145$ $0.0212$ $0.7133$ $0.6264$ $0.7830$
$0.56$ $0.0416$ $0.0368$ $0.0448$ $0.8870$ $0.8179$ $0.9186$
$0.62$ $0.0872$ $0.0818$ $0.0923$ $0.9710$ $0.9617$ $0.9769$
$0.68$ $0.1784$ $0.1726$ $0.1852$ $0.9815$ $0.9775$ $0.9847$
$0.74$ $0.3538$ $0.3462$ $0.3608$ $0.9897$ $0.9879$ $0.9913$
$0.80$ $0.6524$ $0.6407$ $0.6617$ $0.9939$ $0.9934$ $0.9946$
$0.86$ $1.1030$ $1.0812$ $1.1277$ $0.9969$ $0.9964$ $0.9972$
$0.92$ $1.4959$ $1.4892$ $1.5030$ $0.9995$ $0.9994$ $0.9995$
$0.98$ $1.7725$ $1.7665$ $1.7784$ $0.9999$ $0.9999$ $0.9999$
$1.04$ $1.8869$ $1.8780$ $1.8947$ $0.9998$ $0.9997$ $0.9998$
$1.10$ $1.8857$ $1.8754$ $1.8942$ $0.9995$ $0.9994$ $0.9996$
$1.16$ $1.8223$ $1.8110$ $1.8334$ $0.9992$ $0.9990$ $0.9993$
$1.22$ $1.7088$ $1.6951$ $1.7257$ $0.9990$ $0.9986$ $0.9991$
$1.28$ $1.5846$ $1.5678$ $1.6033$ $0.9983$ $0.9979$ $0.9986$
$1.34$ $1.4132$ $1.3864$ $1.4342$ $0.9966$ $0.9942$ $0.9975$
$1.40$ $1.1320$ $1.0944$ $1.1654$ $0.9876$ $0.9832$ $0.9907$
Table 1: Diagnostics summary for the power-law fit for the Glauber dynamics for log-log regression.
Power-law exponents alphaalpha for different sizes $N=512,1024,1536$ are identified for inverse temperature ranges. The log-log regression results are shown in Figure 2a for time power-laws, i.e., $K(t)$ observations over time. This is the functional-diffusion power-laws aligned with the conventional displacement power-laws. We observe anomalous behaviour super-diffusive and sub-diffusive regions, and normal diffusion region being in between. power-law has a different nature as shown in Figure 2b, it doesn’t necessarily represents diffusion behaviour, rather how Monte Carlo dynamics evolve. From simulations perspective this captures the dynamical behaviour due to Glauber spin-flip-dynamics, validating that anomalies present in the dynamics as well. More sophisticated cluster-flips maybe required for slow dynamics regions, such as the Swendsen-Wang [46] dynamics. Uncertainty estimates are achived using bias corrected bootstrapping for reliable confidence interval identification [47, 48]. We demonstrate a set of uncertainty quantification for Glauber dynamics $K(t)$ power-laws log-log plots in Table 1. The similar observations have been made for exponents $\alpha$ the field variations in Figures 3a and 3b. Generalized diffusion coefficients over time validates the same observations, 4a and 4b. Identification of the autocorrelation times from time correlations of average magnetisation. This indicates the different relaxation times matching with diffusion regimes we observed, in Figures 5a and 5b.
df type_power_law N D pvalue parameter
1 Distribution Temperature Range $512$ $0.1875$ $0.9522$ alpha scale
2 Distribution Temperature Range $1024$ $0.1875$ $0.9522$ alpha scale
3 Distribution Temperature Range $1536$ $0.1250$ $0.9998$ alpha scale
4 Distribution Field Range $512$ $0.2500$ $0.7164$ alpha scale
5 Distribution Field Range $1024$ $0.0625$ $1.0000$ alpha scale
6 Distribution Field Range $1536$ $0.1250$ $0.9998$ alpha scale
7 Time Temperature Range $512$ $0.0625$ $1.0000$ alpha scale
8 Time Temperature Range $1024$ $0.0625$ $1.0000$ alpha scale
9 Timer Temperature Range $1536$ $0.1250$ $0.9998$ alpha scale
10 Time Temperature Range $512$ $0.1250$ $0.9998$ C diffusion
11 Time Temperature Range $1024$ $0.1250$ $0.9998$ C diffusion
12 Timer Temperature Range $1536$ $0.1250$ $0.9998$ C diffusion
13 Time Field Range $512$ $0.0625$ $1.0000$ alpha scale
14 Time Field Range $1024$ $0.1875$ $0.9522$ alpha scale
15 Time Field Range $1536$ $0.1250$ $0.9998$ alpha scale
16 Time Field Range $512$ $0.0625$ $1.0000$ C diffusion
17 Time Field Range $1024$ $0.1250$ $0.9998$ C diffusion
18 Time Field Range $1536$ $0.1875$ $0.9522$ C diffusion
Table 2: Diagnostic summary for comparing power-laws on different ranges, with Glauber and Metropolis dynamics via two-sided KS test. Time means $K(t)$ power-laws and distribution means $P(\Gamma(t))$ power-laws.
Finite-size scaling parameters are also computed with nonlinear optimization. We found that $N=1024,1536$ behaves universally. It is also establised that results are statistically significant for both Metropolis and Glauber dynamics 2.
7 Conclusion
A new concept of functional-diffusion is introduced via a canonical example on Ising model i.e., approach to ergodicity for total magnetization. Superdiffusive regimes are identified based on different temperature and field ranges, via computation of power-laws of the forward and inverse fluctuating metric for ergodicity. Quite comprehensive results demonstrate for the first time anomalous convergences to ergodicity quantitatively with power-laws. This concept would be a pedagogical test-bed for the extensions of diffusive behaviour beyond particle-trajectories, so called the functional diffusion for full Ising model and increase the awareness on possible anomalous convergence of metric functionals.
Data Availability
Main diagnostic tables, all data generation and analysis R notebooks are available in public domain at a Zenodo repository [49].
Ackowlegements
We are grateful to Y.Süzen for her kind support of this work. Author would like to express the gratitude to referees for constructive recommendations, making the computational work more robust and pointing out relevant literature.